# American Sign Language Translation ASLP American Sign Language Production

### v3

@@ Abstract
In this paper, we propose a dual-condition diffusion pre-training model named SignDiff that can generate human sign language speakers from a skeleton pose. SignDiff has a novel Frame Reinforcement Network called FR-Net, similar to dense human pose estimation work, which enhances the correspondence between text lexical symbols and sign language dense pose frames, reduces the occurrence of multiple fingers in the diffusion model. In addition, we propose a new method for American Sign Language Production (ASLP), which can generate ASL skeletal pose videos from text input, integrating two new improved modules and a new loss function to improve the accuracy and quality of sign language skeletal posture and enhance the ability of the model to train on large-scale data. We propose the first baseline for ASL production and report the scores of 17.19 and 12.85 on BLEU-4 on the How2Sign dev/test sets. We evaluated our model on the previous mainstream dataset PHOENIX14T, and our method achieved the SOTA results. In addition, our image quality far exceeds all previous results by 10 percentage points in terms of SSIM.
@@ I INTRODUCTION
To accommodate the larger and semantically richer dataset How2Sign, we have developed adaptable modules that optimize mask handling, batch computation, memory utilization, and tensor operations for Fast-SLP Transformers. Additionally, we have designed a novel loss function specifically tailored to the model, which enhances the learning process for longer videos and enables the production of sign language demonstrations of extended durations. Furthermore, we apply SignDiff to create higher-resolution and visually appealing representations of real human signers. SignDiff is the first diffusion model dedicated to producing authentic sign language videos. We developed SignDiff by fine-tuning the widely-used Stable Diffusion 1.5 model [40], training it on carefully processed videos from the How2Sign dataset. Previous GAN-based sign language rendering methods [54, 12, 60, 45, 31] could not meet our demand due to the increasing volume of training data.
@@
  • We propose the first larger-scale American sign language pose video generative model, it demonstrates significant improvements in various capabilities, including the generation of long skeletal pose videos.
@@ II-A Sign Language Production
The development within the field of sign language is not well-balanced, with a concentration of work on Sign Language Recognition (SLR) [13, 20, 28, 25, 10, 29] and Sign Language Translation (SLT) [6, 9, 27, 4]. Various approaches have been explored, including those based on traditional models, neural networks, and model optimization methods. These different techniques have been used in the main research directions. With the development and deepening of the field, there has been increased attention given to Sign Language Production (SLP).
Previously, researchers mainly relied on synthetic animation methods for Sign Language Production (SLP) [11, 26, 50, 36, 35]. However, with the introduction of the Transformer model, many studies have emerged based on neural translation and kinematics. These studies typically focused on PHOENIX14T as a means of exploration and experimentation [44, 46, 23, 47, 48]. While many of these works involve an intermediate step of translating text to gloss, the temporal control of sign language production itself is challenging, leading to potential inaccuracies when translating to gloss.
@@ III-B Dataset Processing
The How2Sign dataset is the first large-scale multimodal and multiview continuous ASL dataset. It includes a parallel corpus comprising over 80 hours of sign language videos and corresponding modalities such as speech, English transcripts, and depth information. Despite its significance, prior researchers in the field of sign language have not effectively employed this corpus due to the intricacies involved in American Sign Language Production (ASLP). Our objective is to train a Sign Language Pose network to generate sequences of 3D skeleton poses on How2Sign as shown in Fig 2. Initially, we extract 2D joint positions from each video using the OpenPose method [7]. We then incorporate the skeletal model estimation improvements introduced in Zelink et al. [70] to transform the 2D joint positions into their corresponding 3D counterparts. Moreover, we employ an iterative inverse kinematics approach to minimize the 3D pose while preserving consistent bone lengths and correcting any misplaced joints. Finally, following a similar approach proposed in Stoll et al. [53], we normalize the skeletal data and adjust the data to a trainable state.
@@ V-A Experimental Setup
After preprocessing, the How2Sign dataset contains 31,047, 1,739, and 2,343 segments respectively. Due to computational limitations and cost considerations, as well as ensuring fair evaluations, we train our models on approximately 8,000 segments from the How2Sign train set. As the first baseline in ASLP, although it is unnecessary to compare our work with previous ones, we also conducted tests using our method on previous mainstream datasets: 1) For experiments that need to be compared to previous work, we run their PHOENIX-14T dataset, such as the back-translation evaluation in the case of Germanic languages. 2) For experiments that do not need to be compared to previous work, we train and test on How2Sign, such as the ablation experiments, multiple evaluations of SignDiff, and Qualitative Evaluation. The training/test data for other models will be mentioned in the Experimental section.
@@ V-B Evaluation for Fast-SLP
Back Translation of ASL. Back translation refers to the process of translating generated videos into spoken sentences, which are then compared with the input sentences. This serves as a popular metric for evaluating SLP [44]. BLEU is a commonly used evaluation metric for this method, with higher values indicating a smoother expression of meaning in the videos. We conducted tests using our new method on the How2Sign dataset and proposed a baseline for the first DL-based ASL video production task, as shown in Table I.
@@ VII CONCLUSION
Providing high-quality sign language to both learning and user communities is of crucial importance for addressing the shortcomings in the community. In this article, we propose the first large-scale production model for ASL videos and the first pose-to-video model based on diffusion models. We introduce a novel FR-Net for the deep integration of sign language and diffusion models. Additionally, we propose two new modules and a new loss function to adapt to our further developed How2Sign ASL dataset on a larger scale. Through baseline comparisons, ablation studies, various parameter analyses, and comprehensive figures, as well as qualitative evaluations, we demonstrate the effectiveness of our approach. In future work, we will enhance the end-to-end capabilities of the model, enabling our technology to be applied to large-scale sign language production.

### v4

@@ Abstract
In this paper, we propose a dual-condition diffusion pre-training model named SignDiff that can generate human sign language speakers from a skeleton pose. SignDiff has a novel Frame Reinforcement Network called FR-Net, similar to dense human pose estimation work, which enhances the correspondence between text lexical symbols and sign language dense pose frames, reduces the occurrence of multiple fingers in the diffusion model. In addition, we propose a new method for American Sign Language Production (ASLP), which can generate ASL skeletal pose videos from text input, integrating two new improved modules and a new loss function to improve the accuracy and quality of sign language skeletal posture and enhance the ability of the model to train on large-scale data. We propose a simple baseline for ASL production and report the scores of 17.19 and 12.85 on BLEU-4 on the How2Sign dev/test sets. We evaluated our model on the previous mainstream dataset PHOENIX14T, and our method achieved the SOTA results. In addition, our image quality far exceeds all previous results by 10 percentage points in terms of SSIM.
@@ I INTRODUCTION
To accommodate the larger and semantically richer dataset How2Sign, we have developed adaptable modules that optimize mask handling, batch computation, memory utilization, and tensor operations for Fast-SLP Transformers. Additionally, we have designed a novel loss function specifically tailored to the model, which enhances the learning process for longer videos and enables the production of sign language demonstrations of extended durations. Furthermore, we apply SignDiff to create higher-resolution and visually appealing representations of real human signers. SignDiff is the a diffusion model dedicated to producing authentic sign language videos (pose2video). We developed SignDiff by fine-tuning the widely-used Stable Diffusion 1.5 model [42], training it on carefully processed videos from the How2Sign dataset. Previous GAN-based sign language rendering methods [56, 12, 62, 47, 33] could not meet our demand due to the increasing volume of training data.
@@
  • We propose a larger-scale American sign language pose video generative model, it demonstrates significant improvements in various capabilities, including the generation of long skeletal pose videos.
@@ II-A Sign Language Production
The development within the field of sign language is not well-balanced, with a concentration of work on Sign Language Recognition (SLR) [13, 22, 30, 27, 10, 31, 17] and Sign Language Translation (SLT) [6, 9, 29, 4]. Various approaches have been explored, including those based on traditional models, neural networks, and model optimization methods. These different techniques have been used in the main research directions. With the development and deepening of the field, there has been increased attention given to Sign Language Production (SLP).
Previously, researchers mainly relied on synthetic animation methods for Sign Language Production (SLP) [11, 28, 52, 38, 37, 18]. However, with the introduction of the Transformer model, many studies have emerged based on neural translation and kinematics. These studies typically focused on PHOENIX14T as a means of exploration and experimentation [46, 48, 25, 49, 50]. While many of these works involve an intermediate step of translating text to gloss, the temporal control of sign language production itself is challenging, leading to potential inaccuracies when translating to gloss.
@@ III-B Dataset Processing
The How2Sign dataset is a large-scale multimodal and multiview continuous ASL dataset. It includes a parallel corpus comprising over 80 hours of sign language videos and corresponding modalities such as speech, English transcripts, and depth information. Despite its significance, prior researchers in the field of sign language have not effectively employed this corpus due to the intricacies involved in American Sign Language Production (ASLP). Our objective is to train a Sign Language Pose network to generate sequences of 3D skeleton poses on How2Sign as shown in Fig 2. Initially, we extract 2D joint positions from each video using the OpenPose method [7]. We then incorporate the skeletal model estimation improvements introduced in Zelink et al. [72] to transform the 2D joint positions into their corresponding 3D counterparts. Moreover, we employ an iterative inverse kinematics approach to minimize the 3D pose while preserving consistent bone lengths and correcting any misplaced joints. Finally, following a similar approach proposed in Stoll et al. [55], we normalize the skeletal data and adjust the data to a trainable state.
@@ V-A Experimental Setup
After preprocessing, the How2Sign dataset contains 31,047, 1,739, and 2,343 segments respectively. Due to computational limitations and cost considerations, as well as ensuring fair evaluations, we train our models on approximately 8,000 segments from the How2Sign train set. As a simple baseline in ASLP, although it is unnecessary to compare our work with previous ones, we also conducted tests using our method on previous mainstream datasets: 1) For experiments that need to be compared to previous work, we run their PHOENIX-14T dataset, such as the back-translation evaluation in the case of Germanic languages. 2) For experiments that do not need to be compared to previous work, we train and test on How2Sign, such as the ablation experiments, multiple evaluations of SignDiff, and Qualitative Evaluation. The training/test data for other models will be mentioned in the Experimental section.
@@ V-B Evaluation for Fast-SLP
Back Translation of ASL. Back translation refers to the process of translating generated videos into spoken sentences, which are then compared with the input sentences. This serves as a popular metric for evaluating SLP [46]. BLEU is a commonly used evaluation metric for this method, with higher values indicating a smoother expression of meaning in the videos. We conducted tests using our new method on the How2Sign dataset and proposed a baseline for the DL-based ASL video production task, as shown in Table I.
@@ VII CONCLUSION
Providing high-quality sign language to both learning and user communities is of crucial importance for addressing the shortcomings in the community. In this article, we propose a large-scale production model for ASL videos and a pose-to-video model based on diffusion models. We introduce a novel FR-Net for the deep integration of sign language and diffusion models. Additionally, we propose two new modules and a new loss function to adapt to our further developed How2Sign ASL dataset on a larger scale. Through baseline comparisons, ablation studies, various parameter analyses, and comprehensive figures, as well as qualitative evaluations, we demonstrate the effectiveness of our approach. In future work, we will enhance the end-to-end capabilities of the model, enabling our technology to be applied to large-scale sign language production.